OpenStudy (anonymous):

How do I determine the vertex of the graph of the function and write the function in standard form x²-6x

OpenStudy (anonymous):

\[ y=x^2- 6 x + 9 -9= (x-3)^2-9 \]

OpenStudy (anonymous):

Complete the square

OpenStudy (anonymous):

x=3 then y=-9

OpenStudy (anonymous):

What do you think the vertex is?